the sum of numbers is 42 but then 6 times the first number exceeds 3 times the second number by 90. find the number
Answer by mananth(16946)

let the numbers be x & y
x+y=42........................1
6x=3y+90
6x-3y=90.......................2
multiply equation 1 by 3
3x+3y=126
add this to equation 2
3x+6x+3y-3y=126+90
9x=216
/9
x=24 one number
other number = 42-24=18
...
m.ananth@hotmail.ca